OpenStudy (anonymous):

Which of the following values for c would mean that the system of equations 2x − 3y = 1 and cx − 3y = 2 would not have any solutions a.2 b.3 c.1 d.0

OpenStudy (imtiaz7):

is your question is complete?

OpenStudy (mathstudent55):

2x - 3y = 1 cx - 3y = 2 Let's take a look at an equation of the form ax + by = c, and see where the slope comes from. Solve for y: \(ax + by = c\) \(by = -ax + c\) \(y = -\dfrac{a}{b}x + \dfrac{c}{b} \) Do you follow this so far?

OpenStudy (mathstudent55):

Since the slope (in the slope-intercept form, or y = mx + b) form is m, our slope is: \(m = -\dfrac{b}{a} \)

OpenStudy (mathstudent55):

As you can see, the slope depends only on a and b, the coefficients of x, and y, respectively. For the system 2x - 3y = 1 cx - 3y = 2 not to have a solution, what must the lines be?

OpenStudy (anonymous):

@mathstudent55 Thank you !!!

OpenStudy (mathstudent55):

What kind of lines have no solution as a system of equations because they never intersect?

OpenStudy (anonymous):

@mathstudent55 parallel ?

OpenStudy (mathstudent55):

Correct. We want the two lines of the given equations to be parallel. What do you know about the slopes of parallel lines?

OpenStudy (anonymous):

@mathstudent55 they have the same slope

OpenStudy (imtiaz7):

yes you are right @kimm210

OpenStudy (mathstudent55):

Correct. Parallel lines have the same slope. Now we know this: Parallel lines have the same slope. The slope of a line depends only on the coefficients of x and y in the equation ax + by = c Since we want the equations to be parallel and have the same slope, that means the coefficients of x and y must be in the same ratio.

OpenStudy (mathstudent55):